OpenStudy (anonymous):

If p equals two, what is the value of three p plus two minus p raised to the second power? Show your work

OpenStudy (anonymous):

3p = 2(3) = 6 6 + 2 = 8 8 - p^2 = 8 - (2)^2 = 8 - 4 = 4 The answer is 4 BAM GURL

OpenStudy (anonymous):

@charlotte123 @kawaiicat123 Mommy I did math :D

OpenStudy (charlotte123):

That Is A Good Boy - I Shall Take You Out For Ice-Cream Today

OpenStudy (anonymous):

LOL CHIBI ^_^

OpenStudy (anonymous):

Where's meh medals c: Who is this Chibi

OpenStudy (charlotte123):

CHI CHI Is Currently Unavailable -

OpenStudy (anonymous):

I AM SURE HE IS CHIBI @_@

OpenStudy (charlotte123):

Nah -

OpenStudy (anonymous):

HE IS

OpenStudy (anonymous):

http://prntscr.com/2bczu9

OpenStudy (charlotte123):

Nope Nope -

OpenStudy (charlotte123):

CONGRATULATIONS EMOTICON!

OpenStudy (anonymous):

whos chibi

OpenStudy (charlotte123):

You Made Your Mother Proud Q_Q

OpenStudy (anonymous):

THAT MOTHER FLAPPIN' CHIBI @_@

OpenStudy (charlotte123):

Nope - CHI CHI - Has Gone Away -

OpenStudy (anonymous):

YOU ARE CHIBI ^_^

OpenStudy (anonymous):

STOP ACTING LIKE A WALNUT @_@

OpenStudy (charlotte123):

Nah Impossible My Dear

OpenStudy (charlotte123):

He he he

OpenStudy (charlotte123):

-Disappears-

OpenStudy (anonymous):

NO CHAR. HE IS CHIBI ;)